# Scaling Laws for Template-Free Detection of Environmental Phase Modulation in Gravitational-Wave Signals

### v1

@@ Abstract
Environmental effects such as hierarchical triple motion introduce cumulative phase modulation in gravitational-wave signals through time-dependent line-of-sight acceleration. Whether such effects are observable depends on both deformation strength and signal-to-noise ratio (SNR). However, this relationship has not been quantified in a waveform-agnostic manner. Here we investigate the detectability of smooth time-warp distortions using template-free time–frequency representations. Rather than relying on direct strain-domain residuals, we analyze trajectory-level statistics derived from the continuous wavelet transform, in particular the evolution of the power-weighted frequency centroid. We show that environmental modulation can be detected using a single-sample statistic referenced to an isolated-binary distribution, without requiring matched templates. Across a grid of cumulative phase distortions $\Delta\phi_{\mathrm{env}}$ and SNR, the detection performance collapses onto a single scaling parameter, $\Lambda=\Delta\phi_{\mathrm{env}}\times\mathrm{SNR}.$ The ROC-AUC follows an approximately sigmoid curve in $\Lambda$ with a clear transition region. For moderate distortions ($\Delta\phi_{\mathrm{env}}\gtrsim 3\,\mathrm{rad}$), environmental modulation is detectable even at low SNR. For smaller distortions ($\Delta\phi_{\mathrm{env}}\sim 1\,\mathrm{rad}$), detectability is noise-limited and emerges only at $\mathrm{SNR}\gtrsim 20$. Within the controlled waveform family studied here, smooth environmental phase modulation is not generically absorbed by intrinsic waveform variability. Instead, detectability is governed by a simple scaling between cumulative phase distortion and signal strength. This scaling relation provides a concrete reference point for template-free environmental searches in gravitational-wave data.
keywords:
gravitational waves, environmental phase modulation, hierarchical triples, time–frequency methods, observability scaling, signal-to-noise ratio
articletype: Paper
1 Introduction
@@
The universe is filled with star systems made up of more than one star. Observation has shown that high mass stars, like O-type, are found predominately in multiple systems [1] and a good chunk of those exist in triples or higher, more so with increasing mass [2]. The systems with three stars are called hierarchical triple systems (HTS) and are important for gravitational wave detection.
@@
Matched filtering using waveform templates representing binary systems have been used successfully for detection and parameter estimation [3]. However, HTS change gravitational waveforms systematically. For HTS motion, the inner binary about a tertiary companion induces a time-dependent line-of-sight acceleration (LOSA) [4, 5] and structured acceleration profiles can produce nontrivial phase evolution beyond constant-acceleration approximations [6, 7]. This can create phase modulation that accumulates over time producing cumulative phase drift. If in detection we assume the source is an isolated binary, but the system is an HTS, this phase drift may be absorbed into intrinsic parameters biasing estimations of mass, chirp rate, or spin. In other words, the LOSA signal stays inside the isolated binary envelope hidden through incorrect parameter estimation – the effect is degenerate.
We therefore ask whether a template-free approach can distinguish isolated (ISO) and LOSA signals. By “template-free” we mean that detectability is evaluated without constructing matched waveform families that explicitly model environmental acceleration. Instead, the analysis relies on distribution-referenced statistics derived from signal representations, rather than direct comparison to environmental templates.
@@
Using time–frequency representations computed from continuous wavelet transforms (CWT) we show that LOSA induces coherent deviations in the trajectory of the power-weighted mean frequency of the signal. We introduce a single-sample, distribution-referenced statistic and evaluate its detectability across a controlled grid space of cumulative phase distortions $\Delta\phi$ and signal-to-noise ratio (SNR). We demonstrate that receiver operating characteristic (ROC) performance approximately collapses onto a one-parameter scaling relation governed by $\Lambda=\Delta\phi_{env}\times\mathrm{SNR}.$ Performance remains near chance below a characteristic $\Lambda$, and rises rapidly above it. In particular, phase distortions $\Delta\phi\gtrsim 3\,\mathrm{rad}$ are detectable even at low SNR, but $\Delta\phi\sim 1\,\mathrm{rad}$ becomes observable only when $\mathrm{SNR}\gtrsim 20$.
@@
This paper shows that LOSA distortions produce measurable shifts in centroid trajectories that follow predictable SNR-$\Delta\phi$ scaling. Their observability is governed by a scaling between cumulative phase shift and SNR. This scaling provides a physically interpretable baseline for template-free environmental searches in current and future GW detectors.
2 Waveform Model and Environmental Phase Modulation
2.1 Isolated Chirp Model
We consider a simplified inspiral-like chirp signal whose strain can be written as
@@ 2.1 Isolated Chirp Model
2.2 Line-of-Sight Acceleration as a Time Warp
@@
2.3 Cumulative Phase Distortion
@@
2.4 Effect on Instantaneous Frequency
@@
To demonstrate that this behavior persists for physically motivated waveforms, we apply the LOSA time reparameterization to a post-Newtonian inspiral generated using the TaylorT4 approximation. Figure 2(a) shows a representative example with cumulative phase distortion $\Delta\phi_{\mathrm{env}}=3\,\mathrm{rad}$ in the noise-free limit. As in the Gaussian case, the deformation manifests primarily as a progressive phase slip rather than as a large amplitude discrepancy. This reflects the fact that LOSA does not alter the intrinsic binary dynamics; instead, it smoothly warps the time coordinate.
For completeness, both figures also display in panel (b) the corresponding deformation in a time–frequency representation. These quantities will be defined formally in Sec. 3.2.
Figure 1: Environmental phase modulation for $\Delta\phi=3\,\mathrm{rad}$ (noise-free). (a) Zoomed strain overlay showing the isolated chirp $h_{\mathrm{iso}}(t)$ (blue) and the LOSA-deformed signal $h_{\mathrm{LOSA}}(t)$ (red) during the final $\sim 0.5\,\mathrm{s}$ prior to peak strain. The dashed curve shows the strain difference $\Delta h(t)=h_{\mathrm{LOSA}}-h_{\mathrm{iso}}$. (b) Time–frequency centroid trajectory and its deviation, which will be defined formally in Sec. 3.2.
Figure 2: Environmental phase modulation applied to a post-Newtonian (TaylorT4) inspiral waveform with $\Delta\phi_{\mathrm{env}}=3\,\mathrm{rad}$ (noise-free). (a) Zoomed strain overlay during the final $\sim 0.4\,\mathrm{s}$ before merger. The LOSA-deformed signal (red) exhibits a cumulative phase slip relative to the isolated waveform (blue); the dashed curve shows the strain difference $\Delta h(t)$. (b) Corresponding time–frequency centroid trajectory $f_{c}(t)$ (top) and its deviation $\Delta f_{c}(t)$ (bottom). The deformation appears as a coherent modification of the centroid evolution, consistent with a smooth time reparameterization rather than a constant phase offset.
@@
3 Experimental Design
The goal of this study is to determine when environmental phase modulation becomes statistically observable in noisy data. We treat detectability as a function of two independent controls: the cumulative environmental phase distortion $\Delta\phi_{\mathrm{env}}$ and the signal-to-noise ratio (SNR). Rather than relying on template matching or reconstruction error, we evaluate a single-sample statistic defined relative to a reference distribution of isolated signals.
3.1 Synthetic Signal Generation
@@
$\Delta\phi_{\mathrm{env}}=\max_{t}\left|\phi_{\mathrm{LOSA}}(t)-\phi_{\mathrm{iso}}(t)\right|.$ (12)
@@
$\Delta\phi_{\mathrm{env}}\in\{0.3,\,1,\,3\}\,\mathrm{rad}\quad\text{and}\quad\mathrm{SNR}\in\{5,\,10,\,20,\,40\}.$
Additional larger distortions are used for diagnostic checks. For each $(\Delta\phi_{\mathrm{env}},\mathrm{SNR})$ pair, isolated and LOSA-modulated samples are generated independently with distinct noise realizations. No pairing or oracle reference is used in the primary statistical evaluation.
3.2 Detection Statistic
@@ 3.3 Scaling Analysis
$\Lambda=\Delta\phi_{\mathrm{env}}\times\mathrm{SNR}.$ (15)
AUROC values from the full $(\Delta\phi_{\mathrm{env}},\mathrm{SNR})$ grid are analyzed as a function of $\Lambda$. This allows us to test for collapse onto a universal curve and to characterize the transition between degenerate and separable regimes via a sigmoid fit.
4 Results
4.1 Detectability Surface in SNR and Cumulative Phase Distortion
Figure 3 shows AUROC as a function of cumulative phase distortion $\Delta\phi_{\mathrm{env}}$ and signal-to-noise ratio (SNR). Each grid point represents an independent evaluation using the centroid-deviation statistic $S_{fc}$.
Figure 3: AUROC as a function of cumulative phase distortion $\Delta\phi_{\mathrm{env}}$ and SNR. Detectability increases with both deformation strength and signal amplitude.
@@ 4.1 Detectability Surface in SNR and Cumulative Phase Distortion
  • For $\Delta\phi_{\mathrm{env}}\approx 0.3$ rad, detectability remains near chance across most SNR values.
  • For $\Delta\phi_{\mathrm{env}}\approx 1$ rad, detection emerges at $\mathrm{SNR}\gtrsim 20$.
  • For $\Delta\phi_{\mathrm{env}}\approx 3$ rad, environmental modulation is detectable even at $\mathrm{SNR}\approx 5$.
@@
4.2 Single-Parameter Scaling and Sigmoid Transition
Figure 4 shows AUROC plotted against $\Lambda$ for all measured grid points. The collapse is striking: data from different SNR and deformation combinations align along a common monotonic trajectory. This indicates that detectability is governed by the product $\Delta\phi_{\mathrm{env}}\times\mathrm{SNR}$, rather than by either parameter independently.
@@ 4.2 Single-Parameter Scaling and Sigmoid Transition
Figure 4: AUROC as a function of the composite parameter $\Lambda=\Delta\phi_{\mathrm{env}}\times\mathrm{SNR}$. Points from all $(\mathrm{SNR},\Delta\phi)$ combinations collapse onto a single curve, demonstrating one-parameter scaling of environmental detectability. The red curve shows a sigmoid fit to the collapsed data. Vertical lines indicate characteristic threshold values corresponding to AUROC = 0.8 and 0.95.
@@
4.3 Time-Local Behavior of the Centroid Shift
@@
for representative deformation strengths. This quantity corresponds to the lower panel (b) in Figs. 1 and 2, where a single example with $\Delta\phi_{\mathrm{env}}=3\,\mathrm{rad}$ was shown. In Fig. 5, we extend that visualization across multiple cumulative phase distortions to examine the scaling behavior of the centroid deviation.
Figure 5: Time evolution of the centroid shift $\Delta f_{c}(t)$ for representative cumulative phase distortions. Larger $\Delta\phi_{\mathrm{env}}$ produces systematically larger centroid deviations, while preserving the characteristic smooth temporal structure.
As the cumulative phase distortion increases, the magnitude of $\Delta f_{c}(t)$ grows smoothly and systematically. The deviation exhibits coherent time structure rather than random fluctuations, confirming that the statistic tracks the expected time-warp deformation of the chirp trajectory. This behavior is consistent with the interpretation developed in Sec. 2: LOSA acts as a smooth reparameterization of time, which manifests as a structured distortion of the time–frequency centroid trajectory.
4.4 Physical Interpretation
The scaling behavior can be understood intuitively. The cumulative phase distortion $\Delta\phi_{\mathrm{env}}$ sets the magnitude of the trajectory deformation, while SNR sets the inverse noise floor against which that deformation is measured. Their product therefore determines the effective signal-to-distortion ratio in the centroid trajectory.
@@ 4.4 Physical Interpretation
5 Discussion
5.1 Physical Interpretation of the Scaling Law
The collapse observed in Sec. 4 suggests $\Lambda$ functions as an effective signal-to-distortion ratio in the centroid trajectory. Across the explored grid of cumulative phase distortions and signal-to-noise ratios, AUROC values collapse onto a single sigmoid curve when plotted against $\Lambda$.
This behavior has a direct physical interpretation. Line-of-sight acceleration induces a smooth time reparameterization, which in turn produces a systematic deviation in the time–frequency trajectory of the signal. In the perturbative regime, the magnitude of this trajectory shift scales approximately linearly with $\Delta\phi_{\mathrm{env}}$. Measurement uncertainty in the centroid trajectory decreases with increasing SNR. Detectability is therefore controlled by the ratio between deterministic deformation amplitude and stochastic measurement noise, naturally leading to the product scaling in Eq. (15).
@@ 5.1 Physical Interpretation of the Scaling Law
5.2 Noise-Limited Versus Physics-Limited Regimes
The detectability surface indicates that weak environmental distortions are not fundamentally hidden by intrinsic waveform structure. Instead, observability is noise-limited at small $\Delta\phi_{\mathrm{env}}$.
For moderate distortions ($\Delta\phi_{\mathrm{env}}\gtrsim 3\,\mathrm{rad}$), detectability remains high even at modest SNR. For smaller distortions ($\Delta\phi_{\mathrm{env}}\sim 1\,\mathrm{rad}$), separation improves rapidly with increasing SNR, becoming reliable only once $\Lambda$ exceeds threshold. Thus the limiting factor in the weak-deformation regime is statistical precision rather than representational degeneracy.
5.3 Practical Use in Detection Pipelines

### v2

# Detectability Scaling Laws for Environmental Phase Modulation in Gravitational-Wave Signals

@@ Abstract
Environmental effects such as hierarchical triple motion introduce cumulative phase modulation in gravitational-wave signals through time-dependent line-of-sight acceleration. Whether such effects are observable depends on both deformation strength and signal-to-noise ratio (SNR). However, this relationship has not been quantified in a waveform-agnostic manner. Here we investigate the detectability of smooth time-warp distortions using template-free time–frequency representations. Rather than relying on direct strain-domain residuals, we analyze trajectory-level statistics derived from the continuous wavelet transform, in particular the evolution of the power-weighted frequency centroid. We show that environmental modulation can be detected using a single-sample statistic referenced to an isolated-binary distribution, without requiring matched templates. Across a grid of cumulative phase distortions $\Delta\phi$ and SNR, the detection performance collapses onto a single scaling parameter, $\Lambda=\Delta\phi\times\mathrm{SNR}.$ The ROC-AUC follows an approximately sigmoid curve in $\Lambda$ with a clear transition region. For moderate distortions ($\Delta\phi\gtrsim 3\,\mathrm{rad}$), environmental modulation is detectable even at low SNR. For smaller distortions ($\Delta\phi\sim 1\,\mathrm{rad}$), detectability is noise-limited and emerges only at $\mathrm{SNR}\gtrsim 20$. Within the controlled waveform family studied here, smooth environmental phase modulation is not generically absorbed by intrinsic waveform variability. Instead, detectability is governed by a simple scaling between cumulative phase distortion and signal strength. This scaling relation provides a concrete reference point for template-free environmental searches in gravitational-wave data.
I Introduction
@@
Many massive stars reside in multiple systems. Observation has shown that high mass stars, like O-type, are found predominately in multiple systems [6] and a good chunk of those exist in triples or higher, more so with increasing mass [7]. The systems with three stars are called hierarchical triple systems (HTS) and are important for gravitational wave detection.
@@
Matched filtering using waveform templates representing binary systems have been used successfully for detection and parameter estimation [3]. However, HTS change gravitational waveforms systematically. Similar phase distortions can arise from other large-scale environmental accelerations, making it useful to understand their detectability in a waveform-agnostic framework. For HTS motion, the inner binary about a tertiary companion induces a time-dependent line-of-sight acceleration (LOSA) [4, 2] and structured acceleration profiles can produce nontrivial phase evolution beyond constant-acceleration approximations [5, 8]. This can create phase modulation that accumulates over time producing cumulative phase drift. If in detection we assume the source is an isolated binary, but the system is an HTS, this phase drift may be absorbed into intrinsic parameters biasing estimations of mass, chirp rate, or spin. In other words, the LOSA signal stays inside the isolated binary envelope hidden through incorrect parameter estimation – the effect is degenerate.
@@
We therefore ask whether environmental phase modulation can be detected without constructing explicit environmental waveform templates. In this work we show that the detectability of such distortions obeys a simple scaling relation governed by the product of cumulative phase deformation and signal-to-noise ratio. In particular, receiver operating characteristic (ROC) performance collapses onto a single parameter $\Lambda=\Delta\phi_{env}\times\mathrm{SNR}.$ Using time–frequency representations computed from continuous wavelet transforms (CWT), we demonstrate this scaling by analyzing trajectory-level statistics derived from the power-weighted frequency centroid of the signal. Detection performance remains near chance below a characteristic value of $\Lambda$, and rises rapidly above it. In particular, phase distortions $\Delta\phi\gtrsim 3\,\mathrm{rad}$ are detectable even at low SNR, but $\Delta\phi\sim 1\,\mathrm{rad}$ becomes observable only when $\mathrm{SNR}\gtrsim 20$.
@@
LOSA distortions produce measurable shifts in centroid trajectories that follow predictable SNR-$\Delta\phi$ scaling. This relation provides a physically interpretable baseline for template-free environmental searches in current and future gravitational-wave detectors.
II Waveform Model and Environmental Phase Modulation
II.1 Isolated Chirp Model
To isolate the time-reparameterization effect under controlled statistical conditions we begin with a minimal inspiral-like chirp model. The same deformation behavior is also demonstrated using a post-Newtonian inspiral waveform.
@@ II.1 Isolated Chirp Model
II.2 Line-of-Sight Acceleration as a Time Warp
@@
II.3 Cumulative Phase Distortion
@@
II.4 Effect on Instantaneous Frequency
@@
Figure 1: Environmental phase modulation for $\Delta\phi=3\,\mathrm{rad}$ (noise-free). (a) Zoomed strain overlay showing the isolated chirp $h_{\mathrm{iso}}(t)$ (blue) and the LOSA-deformed signal $h_{\mathrm{LOSA}}(t)$ (red) during the final $\sim 0.5\,\mathrm{s}$ prior to peak strain. The dashed curve shows the strain difference $\Delta h(t)=h_{\mathrm{LOSA}}-h_{\mathrm{iso}}$. (b) Time–frequency centroid trajectory and its deviation, which will be defined formally in Sec. III.2.
To demonstrate that this behavior persists for physically motivated waveforms, we apply the LOSA time reparameterization to a post-Newtonian inspiral generated using the TaylorT4 approximation. Figure 2(a) shows a representative example with cumulative phase distortion $\Delta\phi=3\,\mathrm{rad}$ in the noise-free limit. As in the Gaussian case, the deformation manifests primarily as a progressive phase slip rather than as a large amplitude discrepancy. This reflects the fact that LOSA does not alter the intrinsic binary dynamics; instead, it smoothly warps the time coordinate.
For completeness, both figures also display in panel (b) the corresponding deformation in a time–frequency representation. These quantities will be defined formally in Sec. III.2.
Figure 2: Environmental phase modulation applied to a post-Newtonian (TaylorT4) inspiral waveform with $\Delta\phi=3\,\mathrm{rad}$ (noise-free). (a) Zoomed strain overlay during the final $\sim 0.4\,\mathrm{s}$ before merger. The LOSA-deformed signal (red) exhibits a cumulative phase slip relative to the isolated waveform (blue); the dashed curve shows the strain difference $\Delta h(t)$. (b) Corresponding time–frequency centroid trajectory $f_{c}(t)$ (top) and its deviation $\Delta f_{c}(t)$ (bottom). The deformation appears as a coherent modification of the centroid evolution, consistent with a smooth time reparameterization rather than a constant phase offset.
@@
III Experimental Design
The goal of this study is to determine when environmental phase modulation becomes statistically observable in noisy data. We treat detectability as a function of two independent controls: the cumulative environmental phase distortion $\Delta\phi$ and the signal-to-noise ratio (SNR). Rather than relying on template matching or reconstruction error, we evaluate a single-sample statistic defined relative to a reference distribution of isolated signals.
III.1 Synthetic Signal Generation
@@
$\Delta\phi=\max_{t}\left|\phi_{\mathrm{LOSA}}(t)-\phi_{\mathrm{iso}}(t)\right|.$ (12)
@@
$\Delta\phi\in\{0.3,\,1,\,3\}\,\mathrm{rad}\quad\text{and}\quad\mathrm{SNR}\in\{5,\,10,\,20,\,40\}.$
Additional larger distortions are used for diagnostic checks. For each $(\Delta\phi,\mathrm{SNR})$ pair, isolated and LOSA-modulated samples are generated independently with distinct noise realizations. No pairing or oracle reference is used in the primary statistical evaluation.
III.2 Detection Statistic
@@ III.3 Scaling Analysis
$\Lambda=\Delta\phi\times\mathrm{SNR}.$ (15)
AUROC values from the full $(\Delta\phi,\mathrm{SNR})$ grid are analyzed as a function of $\Lambda$. This allows us to test for collapse onto a universal curve and to characterize the transition between degenerate and separable regimes via a sigmoid fit.
IV Results
IV.1 Detectability Surface in SNR and Cumulative Phase Distortion
Figure 3 shows AUROC as a function of cumulative phase distortion $\Delta\phi$ and signal-to-noise ratio (SNR). Each grid point represents an independent evaluation using the centroid-deviation statistic $S_{fc}$.
Figure 3: AUROC as a function of cumulative phase distortion $\Delta\phi$ and SNR. Detectability increases with both deformation strength and signal amplitude.
@@ IV.1 Detectability Surface in SNR and Cumulative Phase Distortion
  • For $\Delta\phi\approx 0.3$ rad, detectability remains near chance across most SNR values.
  • For $\Delta\phi\approx 1$ rad, detection emerges at $\mathrm{SNR}\gtrsim 20$.
  • For $\Delta\phi\approx 3$ rad, environmental modulation is detectable even at $\mathrm{SNR}\approx 5$.
@@
IV.2 Single-Parameter Scaling and Sigmoid Transition
Figure 4 shows AUROC plotted against $\Lambda$ for all measured grid points. The collapse is striking: data from different SNR and deformation combinations align along a common monotonic trajectory. This indicates that detectability is governed by the product $\Delta\phi\times\mathrm{SNR}$, rather than by either parameter independently.
@@ IV.2 Single-Parameter Scaling and Sigmoid Transition
Figure 4: AUROC as a function of the composite parameter $\Lambda=\Delta\phi\times\mathrm{SNR}$. Points from all $(\mathrm{SNR},\Delta\phi)$ combinations collapse onto a single curve, demonstrating one-parameter scaling of environmental detectability. The red curve shows a sigmoid fit to the collapsed data. Vertical lines indicate characteristic threshold values corresponding to AUROC = 0.8 and 0.95.
@@
IV.3 Time-Local Behavior of the Centroid Shift
@@
for representative deformation strengths. This quantity corresponds to the lower panel (b) in Figs. 1 and 2, where a single example with $\Delta\phi=3\,\mathrm{rad}$ was shown. In Fig. 5, we extend that visualization across multiple cumulative phase distortions to examine the scaling behavior of the centroid deviation.
Figure 5: Time evolution of the centroid shift $\Delta f_{c}(t)$ for representative cumulative phase distortions. Larger $\Delta\phi$ produces systematically larger centroid deviations, while preserving the characteristic smooth temporal structure.
As the cumulative phase distortion increases, the magnitude of $\Delta f_{c}(t)$ grows smoothly and systematically. The deviation exhibits coherent time structure rather than random fluctuations, confirming that the statistic tracks the expected time-warp deformation of the chirp trajectory. This behavior is consistent with the interpretation developed in Sec. II: LOSA acts as a smooth reparameterization of time, which manifests as a structured distortion of the time–frequency centroid trajectory.
IV.4 Physical Interpretation
The scaling behavior can be understood intuitively. The cumulative phase distortion $\Delta\phi$ sets the magnitude of the trajectory deformation, while SNR sets the inverse noise floor against which that deformation is measured. Their product therefore determines the effective signal-to-distortion ratio in the centroid trajectory.
@@ IV.4 Physical Interpretation
V Discussion
V.1 Physical Interpretation of the Scaling Law
The collapse observed in Sec. IV suggests $\Lambda$ functions as an effective signal-to-distortion ratio in the centroid trajectory. This scaling can be understood using simple phase-measurement arguments. Environmental acceleration produces a deterministic cumulative phase distortion of magnitude $\Delta\phi_{\rm env}$. In gravitational-wave parameter estimation, the uncertainty with which waveform phase can be measured scales approximately as $\sigma_{\phi}\sim 1/\mathrm{SNR}$. Detectability therefore depends on the ratio between deformation amplitude and measurement uncertainty, yielding an effective signal-to-distortion parameter proportional to $\Delta\phi_{\rm env}\times\mathrm{SNR}$. The numerical experiments presented here confirm that ROC performance collapses onto this expected scaling. Across the explored grid of cumulative phase distortions and signal-to-noise ratios, AUROC values collapse onto a single sigmoid curve when plotted against $\Lambda$.
This behavior has a direct physical interpretation. Line-of-sight acceleration induces a smooth time reparameterization, which in turn produces a systematic deviation in the time–frequency trajectory of the signal. In the perturbative regime, the magnitude of this trajectory shift scales approximately linearly with $\Delta\phi$. Measurement uncertainty in the centroid trajectory decreases with increasing SNR. Detectability is therefore controlled by the ratio between deterministic deformation amplitude and stochastic measurement noise, naturally leading to the product scaling in Eq. (15).
@@ V.1 Physical Interpretation of the Scaling Law
The primary contribution of this work is therefore the identification of this simple detectability scaling in $\Lambda=\Delta\phi_{\rm env}\times\mathrm{SNR}$, while the centroid-based statistic used here serves as a representative trajectory diagnostic for demonstrating the effect.
V.2 Noise-Limited Versus Physics-Limited Regimes
The detectability surface indicates that weak environmental distortions are not fundamentally hidden by intrinsic waveform structure. Instead, observability is noise-limited at small $\Delta\phi$.
For moderate distortions ($\Delta\phi\gtrsim 3\,\mathrm{rad}$), detectability remains high even at modest SNR. For smaller distortions ($\Delta\phi\sim 1\,\mathrm{rad}$), separation improves rapidly with increasing SNR, becoming reliable only once $\Lambda$ exceeds threshold. Thus the limiting factor in the weak-deformation regime is statistical precision rather than representational degeneracy.
V.3 Practical Use in Detection Pipelines